# Counting Stacked Objects from Multi-View Images

### v1

@@ Abstract
Figure 1: 3D Counting (3DC). From multiple views of objects to be counted and their container, we estimate both the total volume they occupy and the fraction of this volume taken up by the objects. Combining these two estimates yields the total number of objects.
@@ 1 Introduction
Visual object counting—the task of quantifying the number of instances in a scene—serves as a fundamental building block for numerous real-world applications and autonomous decision-making systems. This challenging computer vision problem spans diverse domains, from cell counting in biomedical imaging [28] and traffic [13] or wildlife [1] monitoring. However, existing counting methods [11, 30, 19, 22, 12] can only count visible objects such as apples spread across a table. The problem becomes significantly harder when objects are stacked in 3D configurations, as in Fig. 1. In such scenarios, occlusion poses a fundamental challenge because only a subset of objects remains visible. This requires models to infer the presence and quantity of hidden instances from limited visual cues. This inference task demands architectures that not only can detect visible object features but also reason about occluded objects through contextual understanding. The challenge is further amplified by variations in stacking patterns, object orientations, and irregular arrangements, making traditional counting approaches inadequate. Accurately counting these stacked objects would address significant industrial and agricultural needs, where precise quantification of items—such as products on a pallet or fruits in crates—not only prevents stock and quality errors but also enhances operational efficiency and logistics.
@@
Figure 2: 3DC pipeline. We decompose the counting task into estimating the volume of the objects to be counted and then estimating the percentage of this volume being used. The first is done on the basis of geometry and second of a depth-map extracted from one of the views.
To this end, we address the task of counting stacked objects from multi-view images. In this setting, while some objects are visible from certain viewpoints, most are occluded beneath the stack. The entire stack, however, can be observed from multiple cameras, allowing for a reasonable estimation of its volume and, importantly, the total object count. We assume the volume of a single object unit is known, which is realistic in scenarios where objects are standardized or belong to a specific class with consistent dimensions, such as packaged goods with minimal variability. We demonstrate that this 3D counting challenge can be solved effectively by decomposing it into two complementary subproblems. As illustrated in Fig. 2, our approach leverages multi-view images of the stack to estimate two key quantities: the 3D geometry of the object stack, and the occupancy ratio of objects within this stack volume. Given the known volume of a single object unit, we compute the total count as the product of the stack volume and occupancy ratio, normalized by the unit volume. This decomposition enables us to solve the 3D counting problem through a combination of geometric reconstruction for volume estimation and deep learning-based depth analysis for occupancy prediction, both of which can be solved efficiently.
@@
We validate our 3D Counting (3DC) approach through extensive experiments on both real-world and synthetic datasets. Our real-world evaluation leverages a diverse collection of scenes depicting diverse objects stacked in containers or still in their original packaging, as illustrated in Fig. 1. To further assess the reliability of 3DC, we also constructed a large-scale synthetic dataset with precisely annotated ground-truth counts. This comprehensive evaluation, along with the public release of our dataset, code, and network weights upon publication, allows for rigorous benchmarking and further advancement of 3D object counting techniques.
@@
  • A complete pipeline for 3D counting of overlapping, stacked objects, a novel and challenging computer vision task that has not been previously addressed in the literature.
  • A network designed to infer the percentage of volume being occupied, which represents a novel idea and forms a key component of the architecture .
  • An extensive new 3D Counting Dataset comprising several thousand physically simulated and rendered scenes with precise ground-truth object counts and volume occupancy computed programmatically.
  • A complementary real-world validation dataset consisting of 20 scenes captured with multi-view images, accurate camera poses, known unit object volumes, and manually verified total counts.
@@ 2 Related work
Counting aims to estimate the number of instances of a specific object category in a scene. Most methods focus on counting visible objects from a single image. A few methods leverage multi-view images to enhance counting accuracy. Our method extends this to a challenging scenario where many of the objects to be counted are occluded, and only multi-view images of the object stack are available. We summarize those related work below.
@@
Most recent counting methods focus on single-view scenarios where they train a network specialized in counting a single object-category such as for crowd-counting [29, 23, 10, 25, 3, 27, 35, 36], counting cars [13], or penguins [1]. These methods are applicable in many domains [2] such as counting cells or other anatomical structures in medical imaging [5] or counting trees or building from satellite images [34]. Single-view counting focuces on addressing challenges like scale variation, perspective distortions, and occlusions. To address these, methods often learn robust feature representations  [39], employ density map estimation [24], or leverage multi-scale features [18]. Apart from the traditional setting that specialized in counting a single category, class-agnostic counting [12, 15] enables counting an arbitrary category at test time, given a few image samples of the class [32], a few bounding boxes [19], or just the class name [30]. Like class-agnostic counting methods, our approach is not restricted to specific categories but can generalize to any object type at test time. For the most part, most counting methods only deal with visible objects. The only related attempt to ours is by Jenkins et al. [6] that infers counts of occluded objects by incoporating LiDAR data into their setup. However, their approach only handles a specific setup of counting different beverages on shelves for a specific set of categories with known volumes, such as “Coca-Cola 20oz bottle” or “milk carton”, while our method generalizes to diverse scenes, object types, and geometries.
@@
Multi-view counting approaches improve accuracy by combining information across multiple camera views, often projecting feature maps onto a common ground plane to generate precise density maps for crowd counting [38, 37] or segmentation maps of fruits [14]. However, these methods assume that all objects are visible from at least one view and are restricted to a specific class of objects, strongly limiting their applicability in real-world scenarios. In contrast, our method targets the largely unexplored area of counting occluded objects, without any restriction on the nature and shapes of these objects.
@@ 3 Method
3DC aims to estimate the total number $\mathcal{N}$ of objects in a container solely from a set of 2D images, which is a challenge even for humans and has not been attempted before to the best of our knowledge.
Our approach is predicated on the idea that, even though we cannot faithfully recover the exact arrangement of all the invisible objects at the bottom of a container, their total number can be inferred by separately estimating their volume in the aggregate, which is a classic vision task, and estimating the object density, which deep learning makes achievable.
After formalizing our approach in Sec. 3.1, we describe our method to performing these two tasks in sections Sec. 3.2 and Sec. 3.3. To train the deep networks we use to estimate density, we created a large synthetic dataset, which we describe in Sec. 3.4 and will make public. Remarkably, as we will see in the results section, even though the networks were trained on this synthetic data, they work effectively on real data as we will see in Sec. 4.
@@ 3.1 Problem Statement
Assuming the average volume occupied by a single object is $v$, and the total volume of the container is $\mathcal{V}$, it would be tempting to compute the number of objects as $\mathcal{N}=\frac{\mathcal{V}}{v}$. However, this fails to account for gaps between objects. If we consider that the objects are stacked in such a way that only a fraction $\gamma$ of the volume $V$ is actually taken up by the objects themselves and that the rest is empty space, then the previous estimate becomes
@@
Our key insight is that this volume usage rate $\gamma$ over the whole container can be estimated with high accuracy from the visible elements only. Note that, in general, the density within the container may not be strictly uniform. However, if it is large enough, the variations tend to compensate each other over the whole volume and using an average value is warranted.
@@
To estimate $\mathcal{N}$ from partial visual information, we assume that:
  • The objects are stacked uniformly.
  • The objects are approximately identical. Their average size $v$ is assumed to be known and does not vary much.
  • Some of the objects are partially visible, so that their density $\gamma$ can be estimated.
These assumptions hold in the real-world scenes of  Fig. 1.
Applications
These assumptions remain general enough for a wide range of applications across various industries. In warehousing and retail, our proposed setup can be used to automate the inventory process by accurately counting stacked items, reducing the need for manual labor and facilitating restocking. Additionally, in manufacturing, our method can enhance quality control by ensuring that shipped containers include a sufficient number of items. It can also provide 3D scene understanding to autonomous systems for robotic tasks like pick-and-place or sorting. Finally, our algorithm can fill an important need in nutrition tracking to estimate the number of items stacked on a plate, allowing for more accurate calorie estimation based on portion size.
@@
Figure 3: Dataset samples. We visualize generated scenes in ascending order of occupied volume ratio, with corresponding depth maps.
@@ 3.2 Volume Estimation
Reconstructing the surface and inferring from it the volume $\mathcal{V}_{est}$ of a set of objects in a container is a well-understood problem. When camera poses are unknown, we use COLMAP [21] to calibrate the images and adjust their scale using a real-world reference measurement. For industrial applications where the camera setup is fixed and known, these calibration and scaling steps are much simpler. Next, we use the SAM2 [20] video segmentation model to isolate the container from its surroundings. When the position of the object stack is unknown, the user only needs to select the stack in the first image, and SAM2 will generate masks for all subsequent images. Given these masks, we extract the container and objects by adding the mask as alpha channel. We then optimize 3D Gaussian Splats [7] from these images. Pixels with alpha lower than 1 are supervised with random pixels each iteration, which in effect prevents gaussians in these pixels and ensures our reconstruction covers the container and objects only.
Finally, to estimate the volume from the reconstruction, we consider the point cloud formed by the means of all gaussians in the 3D reconstruction. We first compute the alpha-concave hull [4] of the points, which defines a mesh that closely envelopes the points while allowing some concavities controlled by a parameter $\alpha$, and thus providing a tighter fit than a convex hull.
Given that the point cloud represents both the container and the objects inside it, this approach effectively isolates the overall shape of both the container and its contents. In a number of cases, the container thickness can be neglected, for example if the stack lies on a flat plane or is inside a thin plastic wrapping. However, if the thickness $T$ of the container cannot be neglected and is specified by the user, we adjust the mesh of the alpha-concave hull to account for the container walls by shifting each vertex inward along its normal by $-T$, thereby refining the estimated volume to represent only the contents. The resulting surface provides us with a good estimate of the volume spanned by the stacked objects.
3.3 Occupied Volume Estimation
Our main intuition is that the volume usage can be very precisely inferred from a depth map computed by a good monocular depth estimator, such as Depth Anything V2 [31] from a view on which the objects of interest are clearly visible. This view is taken to be the camera that is right above the container, unless the objects are better seen from another angle, in which case it can be arbitrarily set by the user. Optionally, this image is then also cropped to cover only the objects of interest.
We therefore aim to learn a function $\Phi:\mathcal{D}\rightarrow\gamma\in[0,1]$ that takes as input a depth image and predicts a volume usage $\gamma$. $\Phi$ learns the relationship between the depth map and the volume usage, including the fact that, if greater depths are visible in the depth map, then the ridges between objects are large and the percentage of volume taken by the object is low. This does not depend on the exact shape of the objects being observed and, should in theory, be applicable to novel inputs with shapes not seen during training. In the results section, we will confirm this to be true in practice.
[…]
Network Architecture.
To implement $\Phi$, we use a DinoV2 [16] encoder model and train a decoder head to predict the percentage of volume used by the stacked objects. This pretrained encoder extracts rich features from the input depth image, encoding meaningful information and making the task of the decoder easier. While it would be possible to generate a very large amount of data and train a specialized model from scratch, we find that leveraging a foundation model drastically improves convergence speed and reduces the need for a large amount of training data. DinoV2 in particular has shown remarkable performance in monocular depth estimation, which motivates our decision to use it for volume usage prediction.
When designing our decoder head, our goal was to aggregate feature values into a single scalar representing the volume usage as predicted from the whole image. To this end, we use consecutive convolutional layers to gradually decrease the resolution of the features, along with ReLU activations, reducing the encoded feature image to a single pixel with 64 channels, and a final linear layer to predict a single scalar from the output of the last convolutional layer. We refer the reader to our supplementary material for additional details on our network architecture.
Training.
We minimize the squared error over the estimated volume usage $\gamma_{est}$. To create ground-truth data for supervision purposes, we physically simulate thousands of objects and measure the exact fraction of volume being used by objects in a container in Sec. 3.4. From these simulated 3D scenes, we generate depth maps with Depth Anything V2 [31], and we augment training images using random resized crops and flips. In an ablation study in Sec. 4, we also report the influence of using these estimated depth maps instead of the perfect ground truth depth maps.
3.4 Dataset generation
@@
in 00000079_RGB0002, 00000079_RGB0013, 00000079_RGB0017

### v2

@@ Abstract
Figure 1: 3D Counting (3DC). From multiple views of objects to be counted and their container, we estimate both the total volume they occupy and the fraction of this volume taken up by the objects. Combining these estimates yields the total number of objects.
@@ 1 Introduction
Visual object counting—the task of quantifying the number of instances in a scene—serves as a fundamental building block for numerous real-world applications and autonomous decision-making systems. This ranges from cell counting in biomedical imaging [27] to traffic [12] and wildlife [1] monitoring. However, these methods [10, 29, 18, 21, 11] can only count visible objects such as apples spread across a table or people in a crowd. The problem becomes significantly harder when objects are stacked on top of each other, as in Fig. 1: Only a subset of them is visible, making counting much more difficult. In fact, our experiments show that this task is truly challenging even for humans. Nevertheless, solving it would have significant applications in industrial and agricultural settings, where precise quantification of items—such as products on a pallet or fruits in crates—not only prevents stock and quality errors but also enhances operational efficiency and logistics.
@@
Figure 2: 3DC pipeline. We decompose the counting task into estimating the volume of the objects to be counted and then estimating the occupancy ratio within that volume. The first is done on the basis of geometry reconstructed from segmentations in multiple images.The second uses as input a depth-map computed by a monocular depth estimator and regresses an occupancy ratio from it.
Overcoming the above-mentioned challenges requires inferring the presence and quantity of hidden instances from limited visual cues. This means not only detecting visible object features but also reasoning about hidden ones through contextual understanding. The challenge is further amplified by variations in stacking patterns, object orientations, and irregular arrangements, making traditional counting approaches inadequate.
At the heart of our proposed solution is a key insight: The number of objects per unit of volume, which we will refer to as the occupancy ratio, can be accurately inferred from a depth map computed by a monocular depth estimator from a view in which enough objects of interest are clearly visible. In most cases, such a view is one where the container is seen roughly from above, without having to be strictly vertical. To exploit it, we break down the problem into two complementary tasks: estimating the 3D geometry of the object stack and estimating the occupancy ratio within this volume, as depicted by Fig. 2. This decomposition enables us to solve the 3D counting problem through a combination of geometric reconstruction for volume estimation and deep learning-based depth analysis for occupancy prediction, both of which can be solved efficiently.
@@
We validate 3DC through extensive experiments on real-world and synthetic datasets. Our real-world evaluation leverages a diverse collection of scenes depicting objects stacked in containers or still in their packaging, as shown in Fig. 1. To further assess the reliability of 3DC, we also constructed a large-scale synthetic dataset with precisely annotated ground-truth counts. This dataset, along with our code, will be made available upon publication.
@@
  • A complete pipeline for 3D counting of overlapping, stacked objects, a novel and challenging computer vision task not previously addressed in the literature.
  • A network designed to infer the occupancy ratio, which embodies a novel idea and forms a key component of the architecture.
  • An extensive new 3D Counting Dataset comprising 400,000 images from 14000 physically simulated and rendered scenes with precise ground-truth object counts and volume occupancy computed programmatically.
  • A complementary real-world validation dataset consisting of 2381 images from 45 scenes captured with accurate camera poses and manually verified total counts.
  • A human baseline derived from 1485 annotations on real images, representing estimates from 33 participants.
In particular, the latter shows that this task is truly hard and that humans perform poorly. This indicates that training a network to predict stacked counts from images in a single step may not be feasible and that our decomposing the problem into simpler subproblems is critical to success, as demonstrated in Sec. 4.
@@ 2 Related work
Counting aims to estimate the number of instances of a specific object category in a scene. Most methods focus on counting visible objects from a single image. A few methods leverage multiple images to enhance counting accuracy. Our method extends this to the more challenging scenario where many of the objects to be counted are hidden. We summarize those related work below.
@@
Most recent counting methods focus on single-view scenarios where they train a network specialized in counting a single object-category such as for crowd-counting [28, 22, 9, 24, 3, 26, 35, 36], counting cars [12], or penguins [1]. These methods are widely applicable [2], such as when counting cells or other anatomical structures in medical imaging [4] or counting trees or building in satellite images [33]. The proposed algorithms address challenges such as scale variation, perspective distortions, and occlusions. Common approaches are to learn robust feature representations  [39], to perform density map estimation [23], or to leverage multi-scale features [17]. Apart from the traditional setting in which objects of a single category are counted, class-agnostic counting [11, 14] enables counting an arbitrary category at test time, given a few image samples of the class [31], a few bounding boxes [18], or just the class name [29]. Like class-agnostic counting methods, our approach is not restricted to specific categories but can generalize to any object type at test time. For the most part, counting methods only deal with visible objects. The only related attempt to ours is the approach of [5] that infers counts of occluded objects by using LiDAR data. However, their algorithm only handles a specific setup of counting different beverages on shelves for a specific set of categories with known volumes, such as “Coca-Cola 20oz bottle” or “milk carton”, while our method does not require LiDAR and generalizes to diverse scenes, object types, and geometries.
@@
Multi-view counting approaches improve accuracy by combining information across multiple camera views, often projecting feature maps onto a common ground plane to generate precise density maps for crowd counting [38, 37] or segmentation maps of fruits [13]. However, these methods also assume that all objects are visible from at least one view and are restricted to a specific class of objects, making them inadequate in many practical scenarios. In contrast, our method targets the largely unexplored area of counting occluded objects, without any restriction on the nature and shapes of these objects.
@@ 3 Method
3DC aims to estimate the total number $\mathcal{N}$ of objects in a container solely from a set of 2D images, which is a challenge even for humans and has not been attempted before as far as we know. Our approach is predicated on the idea that, even though we cannot faithfully recover the exact arrangement of all invisible objects at the bottom of the container, the occupancy ratio can be estimated from a single image, provided that enough objects are visible in it.
To localize the objects in cluttered scenes that may contain several stacks, we rely on a segmentation of the objects $S_{o}$ and their container $S_{c}$ in the first frame. We then use SAM2 [19] to propagate this segmentation as $S_{o,f}$ and $S_{c,f}$ to all subsequent frame $f$, so that the frame in which the most objects are visible can be automatically identified.
In the remainder of this section, we first formalize our approach. We then introduce our occupancy ratio estimator, followed by our approach to estimating volumes.
@@ 3.1 Problem Statement
Assuming the average volume occupied by a single object is $v$, and the total volume of the container is $\mathcal{V}$, it would be tempting to compute the number of objects as $\mathcal{N}=\frac{\mathcal{V}}{v}$. However, this fails to account for gaps between objects. If we consider that the objects are stacked in such a way that only a fraction $\gamma$ of the volume $\mathcal{V}$ is actually taken up by the objects themselves and that the rest is empty space, then the previous estimate becomes
@@
Our key insight is that this volume usage rate $\gamma$ over the whole container can be estimated with high accuracy from the visible elements only. In practice, the density within the container may not be strictly uniform. However, if the container is large enough, the variations tend to compensate each other over the whole volume and using an average value is warranted.
@@
In this work we assume that the objects are stacked uniformly in bulk and approximately identical. Some objects are expected to be partially visible, so that the occupancy ratio $\gamma$ can be estimated.
Applicability.
The above assumptions hold in the real-world scenes of  Fig. 1. Furthermore, they are weak enough to also hold in many realistic scenarios across various industries. In warehousing and retail, our proposed setup can be used to automate the inventory process by accurately counting stacked items, reducing the need for manual labor and easing restocking. In manufacturing, our method can enhance quality control by ensuring that shipped containers include a sufficient number of items. It can also provide 3D scene understanding to autonomous systems for robotic tasks like pick-and-place and sorting.
@@
Figure 3: Dataset samples. We visualize generated scenes in ascending order of occupancy ratio, with ground-truth depth maps.
3.2 Occupancy Ratio Estimation
The most critical step in our approach is estimating the occupancy volume ratio $\gamma$ of Eq. 1 from a single image in which enough target objects can be seen, typically one taken from above even though this is not a strict requirement.
Formally, we seek to learn a function $\Phi:\mathcal{D}\rightarrow\gamma\in[0,1]$ that takes as input a depth map and predicts an occupancy ratio $\gamma$. $\Phi$ learns the relationship between depth maps and occupancy ratio, capturing a key fact: if objects deeper in the stack remain visible, then the gaps between objects are large and the percentage of the volume occupied by them is low. This does not depend on the exact shape of the objects being observed, and should, in theory, apply even to new inputs with shapes not seen during training. In the results section, we will confirm this to be true in practice.
Network Architecture.
To implement $\Phi$, we use an encoder-decoder architecture that first computes rich image features and then decodes them into our target $\gamma$. For the encoder, we use DinoV2 [15], a foundation model trained on many real-world images to help with generalizability while increasing convergence speed.
The decoder has to aggregate feature values into a single scalar representing the volume usage, predicted for the whole image. To this end, we use consecutive convolutional layers to gradually decrease the resolution of the features, along with ReLU activations, reducing the encoded feature image to a single pixel with 64 channels, and a final linear layer to predict a single scalar from the output of the last convolutional layer. We refer the reader to our supplementary material for additional details.
Training Data.
To train $\Phi$, the simplest is to minimize the squared error between the estimated occupancy ratio and ground-truth one over an annotated dataset. Unfortunately, no such dataset exists and we therefore synthesized our own. It comprises 400,000 images spanning 14,000 scenes containing various objects in different containers. Fig. 3 depicts some of them.
To create it, we used the ABC dataset [8] that features a wide variety of computer-assisted design (CAD) models. We retained only watertight objects with a single connected component and rescaled them to fit in a cube of side $0.05$. We then generated a virtual 3D scene with a container, and used a physics-based simulator to drop an initial batch of 100 identical objects in that box. We repeated this step a random number of times or until the box was full, that is to say, the union of objects reached the space above the box after the physical simulation has converged. In each scene, the container is given a random shape and scale. We also include some scenes without any container where objects are directly stacked on the floor, and some scenes where boxes are partially full, as is often the case inreal-life.
Once the physical simulation was complete, we computed analytically the occupancy ratio as well as the total number of objects in the container. To each object and container, we randomly assigned a realistic material that could be metallic, wood, or plastic. In some scenes, we selected a different material for each individual object. Finally, we used a ray-tracing engine to render multiple realistic images of the container and objects seen from several different angles to allow 3D reconstruction, as shown in Fig. 5. Crucially, realistic rendering causes objects at the bottom of the boxes to appear darker due ambient occlusion.
We repeated this for over 14,000 scenes and set aside a subset of 100 to use as a test set of shapes unseen during training. For each one of the 4800 shapes used, we generate a scene with a unit-cube container to reliably measure the ground truth $\gamma$ as well as two additional scenes with random containers. This dataset allows us not only to train $\Phi$ on the top view, but also to run our complete pipeline on the multi-view images in order to measure the accuracy of our count estimate $\mathcal{N}_{est}$, as performed in Sec. 4. We train our model on depth maps produced by the depth estimator, but our dataset also includes ground-truth depth maps, which we employ in an ablation study in our experiments to assess the requirement of accurate depth maps of our method. We report additional statistics that highlight the diversity of the proposed dataset in Fig. 12 of our supplementary.
Inference.
On real-world data, it remains to determine which image to use as input to the $\gamma$-network. We will refer to this image as the key view. To produce the best results, this view should be as close as possible to the depth maps seen during training, as illustraed in  Fig. 4. We automatically select the view that has the largest object segmentation, ensuring the objects are clearly visible, and crop the image to include only the masked content. We then employ Depth Anything V2 [30] to compute a depth map. Note that we also train $\Phi$ on depth maps predicted by that model instead of the ground-truth depth maps, which further reduces the domain gap as demonstrated in our experiments.
(a) (a)
(b) (b)
(c) (c)
Figure 4: Reducing the domain gap. Instead of estimating the occupancy ratio $\gamma$ from synthetic (top) and real images (bottom) (a), we identify a key view (b) and train a network to predict $\gamma$ from their depth maps (c), which are indistinguishable. Top row: synthetic, $\gamma_{gt}=62.4\%$. Bottom row: chocolates, $\gamma_{est}=53.5\%$, $\mathcal{N}_{est}=119$, $\mathcal{N}_{gt}=131$.
@@ 3.3 Volume Estimation
Given an estimate of the occupancy ratio $\gamma$ obtained as discussed above, we still need to compute the total volume of the stack $\mathcal{V}$ to derive the total number of objects from Eq. 1, and the unit volume $v$ if it is not known. When the cameras are uncalibrated, we use COLMAP [20] to compute their poses and adjust their scale using a real-world reference measurement. In industrial scenarios with fixed camera setups known a priori, simpler methods could be be used.
Inferring the volume $\mathcal{V}_{est}$ of the stacked objects from multiple images is a well-understood problem. In our specific implementation, we start from the segmented images and extract the container and objects by adding the mask as alpha channel. We then optimize 3D Gaussian Splats [6] from these images, ensuring our reconstruction covers the container and objects only. To compute the volume the splats enclose, we adapt the voxel carving algorithm. We initialize a voxel grid from the bounding box of the gaussians. Then, given the masks generated previously and depth maps rendered from 3DGS, a voxel is carved out if its projection on any given view falls outside of the mask or if its projected depth is less than the reconstructed depth map. This procedure successfully reconstructs objects and their containers, including partially filled boxes.
@@
in 4_0, 4_1, 4_2
Figure 5: Multi-view images. We generate 30 views from arbitrary angles for each of the scenes in our large-scale synthetic dataset.
To remove the container from the reconstruction, we estimate its thickness $t$ and erode the voxels on all sides except the top by $t$, thereby refining the estimated volume to represent only the contents. The value of $t$ is predicted by an additional decoder $\Psi$ that takes the same encoded image features as our previous network $\Phi$ in Sec. 3.2. To this end, we use dilated convolutional layers [34], increasing the network’s receptive field with no additional parameters. To make this prediction scale independent and easily predictable from 2D images alone, we predict thickness as a ratio of the container’s size. We supervise $\Psi$ with ground-truth thickness from our dataset and at inference we average the estimation of $\Psi$ over all images. Taken together, these algorithms yield a good estimate of the volume $\mathcal{V}$ spanned by the stacked objects.
Unit volume $v$.
In most applications, the unit volume $v$ of Eq. 1 is known exactly because the object has been manufactured to a precise specification, $v$ can be computed from the simple geometry of the objects, or $v$ can be obtained from existing reference data, particularly for food items. When the unit volume of an object $v$ is not readily available, we estimate its value using the method described above and from a set of images of a template object. This task is made easier by the absence of a container, and this is shape but not scene specific. For a new shape, one exemplar is selected in a single frame and we can then use SAM2 [19] to generate a segmentation on all frames at once. The unit volume $v$ computed this way can then be used across all scenes containing this object.
4 Experiments
We evaluate our method in two ways: measuring the accuracy of 3D counting as a whole and of the occupancy ratio estimation, in Sec. 4.2 and Sec. 4.3, respectively. These evaluations are performed over two datasets. The first comprises 100 scenes representing physically simulated shapes from the ABC dataset [8]. These scenes were isolated after their generation and were not seen during the training of our occupancy ratio network. The second is made of 2381 real images spanning 45 real scenes that were captured with a regular smartphone’s RGB camera, and no additional sensor. These captures offer multiple views around stacks of objects in a container, lying flat on a table or still enclosed in their packaging. We count the ground-truth number of units manually for all scenes below 500, or infer it from the weight for even larger counts. In Fig. 8, we provide intermediate results to help our readers form a better intuition about our method’s behavior, and additional qualitative results in Fig. 6 and Fig. 1.